OpenStudy (anonymous):

Evaluate (if possible) the six trigonometric functions of the real number. (If not possible, enter IMPOSSIBLE.) t = 5π/6

OpenStudy (anonymous):

sint = 1/2 cost = -sqrt(3)/2 tant = -1/sqrt(3) csct = 2 sect = -2/sqrt(3) cott = -sqrt(3)

OpenStudy (anonymous):

work out sine and cos from graph, then tan = sin/cos csc = 1/sin sec = 1/cos cot = 1/tan

OpenStudy (anonymous):

cost = -sqrt(3)/2 is this \[\sqrt{-3}/2\]

OpenStudy (anonymous):

\[\cos(t) = -\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

the negative is outside

OpenStudy (anonymous):

if you remember how to convert radians to degrees, then know the standard values for 0, 30 , 60 , 45, 90 , 180 its really useful.